1. 1.275 liter of saturated silver chloride AgCl contains 0.01892 g of

dissolved AgCl at 25℃. Calculate the Ksp for AgCl.

2/ Balance the redox reaction in a) acidic and b) basic solution.

Fe2+ + Cr2O7 2- -> Fe3+ + Cr3+

3 A sample containing iron is titrated with a 0.025 M potassium

dichromate K2Cr2O7 solution with H3PO4/H2SO4. The titration required

26.74 mL of K2Cr2O7 for 1.45 g sample. What is the percent iron in the

sample? (Use the balanced reaction in #2).


1
Expert's answer
2022-06-06T09:06:03-0400

Solution (1):

The molar mass of AgCl is 143.32 g/mol

Therefore,

Moles of AgCl = (0.001892 g AgCl) × (1 mol AgCl / 143.32 g AgCl) = 0.0000132 mol AgCl


Molarity = Moles of solute / Liters of solution

Therefore,

Molarity of AgCl = (0.0000132 mol) / (1.275 L) = 1.035×10−5 mol/L

Molarity of AgCl = 1.035×10−5 M


AgCl(s) → Ag+(aq) + Cl(aq)

According to stoichiometry:

[Ag+] = [Cl] = Molarity of AgCl = 1.035×10−5 M


The Ksp expression for AgCl:

Ksp = [Ag+] × [Cl]

Therefore,

Ksp for AgCl = (1.035×10−5) × (1.035×10−5) = 1.07×10−10

The Ksp for AgCl is 1.07×10−10



Solution (2):

Fe2+ + Cr2O72− → Fe3+ + Cr3+

a) in acidic solution:

Oxidation half-reaction:

Fe2+ − e− → Fe3+

Reduction half-reaction:

Cr2O72− + 14H+ + 6e− → 2Cr3+ + 7H2O

Balance the redox reaction in acidic solution:

6Fe2+ + Cr2O72− + 14H+ → 6Fe3+ + 2Cr3+ + 7H2O


b) in basic solution Cr2O72− is converted to CrO42−:

Cr2O72− + 2OH → 2CrO42− + H2O

Oxidation half-reaction:

Fe2+ − e− → Fe3+

Reduction half-reaction:

CrO42− + 4H2O + 3e→ Cr(OH)3 + 5OH

Balance the redox reaction in basic solution:

3Fe2+ + CrO42− + 4H2O → 3Fe3+ + Cr(OH)3 + 5OH



Solution (3):

Balance the redox reaction in acidic solution:

6Fe2+ + Cr2O72− + 14H+ → 6Fe3+ + 2Cr3+ + 7H2O

According to stoichiometry:

Moles of Fe2+ / 6 = Moles of Cr2O72−

or:

Moles of Fe2+ = 6 × Molarity of Cr2O72− × Volume of Cr2O72−

Moles of Fe2+ = 6 × 0.025 M × 0.02674 L = 0.0040 mol Fe2+


Fe → Fe2+

According to stoichiometry:

Moles of Fe = Moles of Fe2+ = 0.0040 mol


The molar mass of iron (Fe) is 55.845 g/mol

Therefore,

Mass of Fe = (0.0040 mol Fe) × (55.845 g Fe / 1 mol Fe) = 0.22338 g Fe


%Fe in the sample = (Mass of Fe / Mass of sample) × 100%

%Fe in the sample = (0.22338 g / 1.45 g) × 100% = 15.4%

The percent iron (Fe) in the sample is 15.4%